As a prospective share holder, you are offered a project that will pay you an income of #8000 in 5years time at 13% rate of discount. What is the present value your future dividend.
1
Expert's answer
2021-08-03T13:28:51-0400

Solution:

Given, "FV=8000,i=13\%=0.13,n=5"

"i+1=(\dfrac{FV}{PV})^{1/n}"

Putting given values:

"0.13+1=(\dfrac{8000}{PV})^{1/5}
\\\Rightarrow 1.13^5=\dfrac{8000}{PV}
\\\Rightarrow PV=4342.79"